OpenStudy (anonymous):

The first term of a geometric series is 3, and the sum of the first term and the second term is 15. What is the sum of the first six terms? 1,023 3,906 4,096 11,718

OpenStudy (anonymous):

@bibby @TheSaint905621

OpenStudy (anonymous):

@ilikemath50

OpenStudy (paxpolaris):

the General term of geometric sequence is \[\large a_n=a \cdot r^{n-1}\] \[a=3\]\[a+ar=15\] what's r?

OpenStudy (anonymous):

12?

OpenStudy (paxpolaris):

\[3+3r=15 \]\[3r=12\]\[r=?\]

OpenStudy (anonymous):

idk? :(

OpenStudy (anonymous):

@ilovebmth1234

OpenStudy (anonymous):

@PaxPolaris you still there?

OpenStudy (paxpolaris):

r =4 now that we have both first term and common ratio we can find the sum of the first 6 terms...

OpenStudy (paxpolaris):

\[\large {S_6=a+ar+ar^2+ar^3+ar^4+ar^5\\ \ \ \ \ \ = a \left( 1+r+r^2+r^3+r^4+r^5 \right)}\]...

OpenStudy (paxpolaris):

\[\Large \therefore S_6=\color{green}{a \cdot {r^6-1\over r-1}}\]

OpenStudy (paxpolaris):

? none of the choices seems right ?

OpenStudy (anonymous):

agreed but one of them is the answer and I have no idea what it is

OpenStudy (ilovebmth1234):

the answer is 4096

OpenStudy (paxpolaris):

the answer is\[\large \cancel3\cdot {4^6-1 \over \cancel{4-1}}=4096 -1 = 4095\] it's not 4096

OpenStudy (ilovebmth1234):

yeah ik i got 4095 but she s my sister and she put the 4096 and it was right i did it all on paper infront of her so both of us are right ha

OpenStudy (anonymous):

4096 was right in plato